# Ref-GS: Directional Factorization for 2D Gaussian Splatting

### v1

@@ 1 Introduction
This is because NeRF and its follow-ups [2, 3, 38, 45] represent 3D scenes as a collection of emission radiance points and query view-dependent colors using viewing direction, without accounting for the bouncing and bending of light rays as they travel from the light source to the viewing cameras. To tackle this issue, Ref-NeRF [36] take advantage of surface light field rendering [40, 5] and replaces NeRF’s directional parameterization with an integrated reflection encoding, achieving significant improvement in the realism and accuracy of specular reflections. Recent work [41] brings feature grid-based encoding to the directional domain to speed up the efficiency of directional encoding. In the context of 3DGS, directly applying the reflection of the view direction as the view-dependent color query for recent efficient GS representation is problematic, as it independently inherits model orientation and Spherical Harmonics (SH) color for each primitive, resulting in transforming viewing direction can easily be offset during parameter updating. To do so, recent work [43, 46] incorporates smooth regularization and higher-order view-dependent color modeling into the rendering function, achieving promising quality on reflective surfaces. Despite the high rendering quality, they are struggling to provide accurate geometry.
@@ 4 Ambiguity in Directional Query
In prior Gaussian Splatting methods, the diffusion, reflection and refraction components of each primitive are simplified using view-dependent emission radiance, which significantly accelerates the forward rendering process without the needed of per instant lighting evaluation. Then, they optimize the emission radiance together with the geometry jointly through an inverse rendering framework by backpropagating a multi-view photometric loss.
@@ 5.3 Far-field Lighting
We now present a novel $\operatorname{Sph-Mip}$ encoding for modeling high-frequency far-field lighting, using a learnale multi-level spherical feature grid, named $\operatorname{Sph-Mip}$ grid.
@@
Note that, as shown in Fig. 2, the $\operatorname{Sph-Mip}$ grid is three-dimensional, the directional coordinates $(\theta,\phi)$ correspond to the XY axes of the grid, while the Z-axis represents the roughness $\rho$ variance. Given reasterized buffers, we first calculate its corresponding spherical coordinates $(\theta,\phi)$ for each pixel by:
@@ A.1 Network
The goal of the shallow MLP $f_{\Theta}$ is is to non-linearly map the directional feature $\mathbf{S}\in\mathbb{R}^{H\times W\times 16}$ produced by the $\operatorname{Sph-Mip}$ encoding and the high-dimensional intermediate tensor $\mathbf{K}\otimes\mathbf{S}$ has a shape of $H\times W\times 64$. Our MLP accepts an input having $16+64$ feature dimensions. The input is fed into a 2-layer MLP with 256 neurons per hidden layer in them followed by ReLU activation functions. The output is fed into a output head predicts the view-dependent radiance with a exponential function output layer. Finally, we apply gamma tone mapping [1] $\gamma(\cdot)$ to convert the colors into the sRGB space before calculating the rendering loss:
@@ C.4 NeRF Synthetic Dataset
C.5 Additional Results on Real-World Captures
@@
C.6 Scene Decompositions and Editing
@@
C.7 Supplementary Video Results

### v2

@@ 1 Introduction
This is because NeRF and its follow-ups [3, 4, 37, 45] represent 3D scenes as a collection of emission radiance points and query view-dependent colors using viewing direction, without accounting for the bouncing and bending of light rays as they travel from the light source to the viewing cameras. To tackle this issue, Ref-NeRF [35] take advantage of surface light field rendering [40, 6] and replaces NeRF’s directional parameterization with an integrated reflection encoding, achieving significant improvement in the realism and accuracy of specular reflections. Recent work [41] brings feature grid-based encoding to the directional domain to speed up the efficiency of directional encoding. In the context of 3DGS, directly applying the reflection of the view direction as the view-dependent color query for recent efficient GS representation is problematic, as it independently inherits model orientation and Spherical Harmonics (SH) color for each primitive, resulting in transforming viewing direction can easily be offset during parameter updating. To do so, recent work [43, 46, 39] incorporates smooth regularization and higher-order view-dependent color modeling into the rendering function, achieving promising quality on reflective surfaces. Despite the high rendering quality, they are struggling to provide accurate geometry.
@@ 4 Ambiguity in Directional Query
In prior Gaussian Splatting methods, the diffusion, reflection and refraction components of each primitive are simplified using view-dependent emission radiance, which significantly accelerates the forward rendering process without the need for per-instance lighting evaluation. Then, they optimize the emission radiance together with the geometry jointly through an inverse rendering framework by backpropagating a multi-view photometric loss.
@@ 5.3 Far-field Lighting
We now present a novel $\operatorname{Sph-Mip}$ encoding for modeling high-frequency far-field lighting, using a learnable multi-level spherical feature grid, named $\operatorname{Sph-Mip}$ grid.
@@
Note that, as shown in Fig. 2, the $\operatorname{Sph-Mip}$ grid is three-dimensional, the directional coordinates $(\theta,\phi)$ correspond to the XY axes of the grid, while the Z-axis represents the roughness $\rho$ variance. Given rasterized buffers, we first calculate its corresponding spherical coordinates $(\theta,\phi)$ for each pixel by:
@@ 7 Conclusion
Acknowledgements
This work is supported by the National Natural Science Foundation of China (NSFC No.62272184) and “Pioneer” and “Leading Goose” R&D Program of Zhejiang (No.2024C01161). The computation is completed in the HPC Platform of Huazhong University of Science and Technology.
@@ A.1 Network
The goal of the shallow MLP $f_{\Theta}$ is is to non-linearly map the directional feature $\mathbf{S}\in\mathbb{R}^{H\times W\times 16}$ produced by the $\operatorname{Sph-Mip}$ encoding and the high-dimensional intermediate tensor $\mathbf{K}\otimes\mathbf{S}$ has a shape of $H\times W\times 64$. Our MLP accepts an input having $16+64$ feature dimensions. The input is fed into a 2-layer MLP with 256 neurons per hidden layer in them followed by ReLU [1] activation functions. The output is fed into a output head predicts the view-dependent radiance with a exponential function output layer. Finally, we apply gamma tone mapping [2] $\gamma(\cdot)$ to convert the colors into the sRGB space before calculating the rendering loss:
@@ C.4 NeRF Synthetic Dataset
C.5 Additional Ablation Results
We provide more ablation results of on synthesized test in Tab. 8. To more clearly demonstrate the distinct advantages of the 2D Gaussian representation, we replaced 2DGS [11] with 3DGS [14], using the shortest axis as the plane normal while keeping the rest unchanged for comparison, as shown in the first two rows of Tab. 8. Furthermore, We have conducted ablation studies on the grid size $N$ of $\operatorname{Sph-Mip}$, as shown in Tab. 8. Notably, 3DGS-DR[46] improves the performance of GaussianShader[12] by introducing deferred shading with a simple shading model. “w/o $\mathbf{K}\otimes\mathbf{S}$” demonstrates that the $\operatorname{Sph-Mip}$ encoding can further enhance rendering quality. Additionally, the results of “w/o DS” demonstrate that our method outperforms the explicit BRDF of GaussianShader.
Chair Drums Lego Mic Materials Ship Hotdog Ficus
Ours 34.66 26.33 36.26 35.76 30.99 29.67 37.39 34.52
w/ 3DGS 34.15 25.86 34.74 34.73 31.32 29.52 36.78 33.10
$\operatorname{Sph-Mip}$ $N$=8 34.67 26.34 35.83 35.23 30.91 29.26 37.19 34.11
$\operatorname{Sph-Mip}$ $N$=7 34.64 26.36 35.80 35.17 31.00 29.23 37.10 34.10
$\operatorname{Sph-Mip}$ $N$=6 35.65 26.17 35.74 35.04 30.39 29.16 37.13 34.10
w/o DS 33.75 25.85 33.99 35.16 29.25 28.89 36.11 32.15
w/o $\mathbf{K}\otimes\mathbf{S}$ 34.08 25.71 35.19 34.21 29.77 29.10 36.62 32.47
Table 8: Per-scene PSNR comparison on NeRF Synthetic dataset. w/ 3DGS: Using 3DGS as the representation of our Ref-GS with the rest unchanged.
C.6 Additional Results on Real-World Captures
@@
C.7 Scene Decompositions and Editing
@@
C.8 Supplementary Video Results